# Unexpected Improvements to Expected Improvement for Bayesian Optimization

### v2

@@ 1 Introduction
Bayesian Optimization (BO) is a widely used and effective approach for sample-efficient optimization of expensive-to-evaluate black-box functions (Frazier, 2018; Garnett, 2023), with applications ranging widely between aerospace engineering Lam et al. (2018), biology and medicine Langer and Tirrell (2004), materials science Ament et al. (2021), civil engineering Ament et al. (2023), and machine learning hyperparameter optimization Snoek et al. (2012); Turner et al. (2021). BO leverages a probabilistic surrogate model in conjunction with an acquisition function to determine where to query the underlying objective function. Improvement-based acquisition functions, such as Expected Improvement (EI) and Probability of Improvement (PI), are among the earliest and most widely used acquisition functions for efficient global optimization of non-convex functions (Močkus, 1975; Jones et al., 1998). EI has been extended to the constrained (Gelbart et al., 2014; Gardner et al., 2014), noisy (Letham et al., 2019), and multi-objective (Emmerich et al., 2006) setting, as well as their respective batch variants (Wilson et al., 2018; Balandat et al., 2020; Daulton et al., 2020), and is a standard baseline in the BO literature (Snoek et al., 2012; Frazier, 2018). While much of the literature has focused on developing new sophisticated acquisition functions, subtle yet critical implementation details of foundational BO methods are often overlooked. Importantly, the performance of EI and its variants is inconsistent even for mathematically identical formulations and, as we show in this work, most often sub-optimal.
Although the problem of optimizing EI effectively has been discussed in various works, e.g. (Frazier, 2018; Wilson et al., 2018; Gramacy et al., 2022), prior focus has been on optimization algorithms and initialization strategies, rather than the fundamental issue of computing EI.
@@
    We extend the ideas behind analytical LogEI to other members of the EI family, including constrained EI (CEI), Expected Hypervolume Improvement (EHVI), as well as their respective batch variants for parallel BO, qEI and qEHVI, using smooth approximations of the acquisition utilities to obtain non-vanishing gradients. All of our methods are available as part of BoTorch (Balandat et al., 2020).
@@
To increase the chance of finding the global optimum of non-convex functions, gradient-based optimization is typically performed from multiple starting points, which can help avoid getting stuck in local optima (Törn and Zilinskas, 1989). For improvement-based acquisition functions however, optimization becomes increasingly challenging as more data is collected and the likelihood of improving over the incumbent diminishes, see our theoretical results in Section 3 and the empirical illustration in Figure 1 and Appendix D.2. As a result, gradient-based optimization with multiple random starting points will eventually degenerate into random search when the gradients at the starting points are numerically zero. This problem is particularly acute in high dimensions and for objectives with a large range.
@@ 2.1 Gaussian Processes
Gaussian Processes (GP) (Rasmussen, 2004) are the most widely used surrogates in BO, due to their high data efficiency and good uncertainty quantification. For our purposes, it suffices to consider a GP as a mapping that provides a multivariate Normal distribution over the outputs $f({\mathbf{x}})$ for any ${\mathbf{x}}$:
@@ 2.2 Improvement-based Acquisition Functions
For the fully-sequential ($q=1$), single-outcome ($M=1$) setting, “classic” EI (Mockus, 1978) is defined as
@@
Constrained BO involves one or more black-box constraints and is typically formulated as finding $\max_{{\mathbf{x}}\in\mathbb{X}}f_{\textrm{true},1}({\mathbf{x}})$ such that $f_{\textrm{true},i}({\mathbf{x}})\leq 0$ for $i\in\{2,\dotsc,M\}$. Feasibility-weighting the improvement (Gardner et al., 2014; Gelbart et al., 2014) is a natural approach for this class of problems:
@@
In many settings, one may evaluate $f_{\textrm{true}}$ on $q>1$ candidates in parallel to increase throughput. The associated parallel or batch analogue of EI (Ginsbourger et al., 2008; Wang et al., 2016) is given by
@@
Unlike EI, qEI does not admit a closed-form expression and is thus typically computed via Monte Carlo sampling, which also extends to non-Gaussian posterior distributions (Wang et al., 2016; Balandat et al., 2020):
@@
due to observing a batch ${\mathbf{f}}({\mathbf{X}}):=[{\mathbf{f}}({\mathbf{x}}_{1}),\cdots,{\mathbf{f}}({\mathbf{x}}_{q})]$ of $q$ new observations. EHVI can be expressed in closed form if $q=1$ and the objectives are modeled with independent GPs (Yang et al., 2019), but Monte Carlo approximations are required for the general case (qEHVI) (Daulton et al., 2020).
@@ 2.3 Optimizing Acquisition Functions
Optimizing an acquisition function (AF) is a challenging task that amounts to solving a non-convex optimization problem, to which multiple approaches and heuristics have been applied. These include gradient-free methods such as divided rectangles (Jones et al., 1993), evolutionary methods such as CMA-ES (Hansen et al., 2003), first-order methods such as stochastic gradient ascent, see e.g., Wang et al. (2016); Daulton et al. (2022b), and (quasi-)second order methods (Frazier, 2018) such as L-BFGS-B (Byrd et al., 1995). Multi-start optimization is commonly employed with gradient-based methods to mitigate the risk of getting stuck in local minima. Initial points for optimization are selected via various heuristics with different levels of complexity, ranging from simple uniform random selection to BoTorch’s initialization heuristic, which selects initial points by performing Boltzmann sampling on a set of random points according to their acquisition function value (Balandat et al., 2020). See Appendix B for a more complete account of initialization strategies and optimization procedures used by popular implementations. We focus on gradient-based optimization as often leveraging gradients results in faster and more performant optimization (Daulton et al., 2020).
Optimizing AFs for parallel BO that quantify the value of a batch of $q>1$ points is more challenging than optimizing their sequential counterparts due to the higher dimensionality of the optimization problem – $qd$ instead of $d$ – and the more challenging optimization surface. A common approach to simplify the problem is to use a sequential greedy strategy that greedily solves a sequence of single point selection problems. For $i=1,\dotsc,q$, candidate ${\mathbf{x}}_{i}$ is selected by optimizing the AF for $q=1$, conditional on the previously selected designs $\{{\mathbf{x}}_{1},...,{\mathbf{x}}_{i-1}\}$ and their unknown observations, e.g. by fantasizing the values at those designs (Wilson et al., 2018). For submodular AFs, including EI, PI, and EHVI, a sequential greedy strategy will attain a regret within a factor of $1/e$ compared to the joint optimum, and previous works have found that sequential greedy optimization yields improved BO performance compared to joint optimization (Wilson et al., 2018; Daulton et al., 2020). Herein, we find that our reformulations enable joint batch optimization to be competitive with the sequential greedy strategy, especially for larger batches.
@@ 2.4 Related Work
While there is a substantial body of work introducing a large variety of different AFs, much less focus has been on the question of how to effectively implement and optimize these AFs. Zhan and Xing (2020) provide a comprehensive review of a large number of different variants of the EI family, but do not discuss any numerical or optimization challenges. Zhao et al. (2023) propose combining a variety of different initialization strategies to select initial conditions for optimization of acquisition functions and show empirically that this improves optimization performance. However, they do not address any potential issues or degeneracies with the acquisition functions themselves. Recent works have considered effective gradient-based approaches for acquisition optimization. Wilson et al. (2018) demonstrates how stochastic first-order methods can be leveraged for optimizing Monte Carlo acquisition functions. Balandat et al. (2020) build on this work and put forth sample average approximations for MC acquisition functions that admit gradient-based optimization using deterministic higher-order optimizers such as L-BFGS-B.
Another line of work proposes to switch from BO to local optimization based on some stopping criterion to achieve faster local convergence, using either zeroth order (Mohammadi et al., 2015) or gradient-based (McLeod et al., 2018) optimization. While McLeod et al. (2018) are also concerned with numerical issues, we emphasize that those issues arise due to ill-conditioned covariance matrices and are orthogonal to the numerical pathologies of improvement-based acquisition functions.
@@ 3 Theoretical Analysis of Expected Improvement’s Vanishing Gradients
For any given – and especially early – iteration, $\epsilon_{n}$ does not have to be small, as both the optimality gap and the maximal posterior standard deviation can be large initially. Note that under certain technical conditions on the kernel function and the asymptotic distribution of the training data $\mathcal{D}_{n}$, the maximum posterior variance vanishes guaranteeably as $n$ increases, see (Lederer et al., 2019, Corollary 3.2). On its own, Theorem 1 gives insight into the non-asymptotic behavior by exposing a dependence to the distribution of objective values $f$. In particular, if the set of inputs that give rise to high objective values ($\approx f^{*}$) is concentrated, $P(f({\mathbf{x}})<f^{*}-\epsilon)$ will decay very slowly as $\epsilon$ increases, thereby maintaining a lower bound on the probability of close to 1. As an example, this is the case for the Ackley function, especially as the dimensionality increases, which explains the behavior in Figure 1.
@@ 4.1 Analytic LogEI
where $c_{1}=\log(2\pi)/2$, and $c_{2}=\log(\pi/2)/2$, $\epsilon$ is the numerical precision, and log1mexp, erfcx are numerically stable implementations of $\log(1-\exp(z))$ and $\exp(z^{2})\text{erfc}(z)$, respectively, see App. A. Progenitors of Eq. (9) are found in SMAC 1.0 Hutter et al. (2011) and RoBO Klein et al. (2017) which contain a log-transformed analytic EI implementation that is much improved, but can still exhibit instabilities as $z$ grows negative, see App. Fig. 10. To remedy similar instabilities, we put forth the third, asymptotic case in Eq. (9), ensuring numerical stability throughout, see App. A.2 for details. The asymptotically quadratic behavior of log_h becomes apparent in the last two cases, making the function particularly amenable to gradient-based optimization with significant practical implications for EI-based BO.
@@ 4.4 Monte Carlo Parallel LogEHVI
The numerical difficulties of qEHVI in (6) are similar to those of qEI, and the basic ingredients of smoothing and log-transformations still apply, but the details are significantly more complex since qEHVI uses many operations that have mathematically zero gradients with respect to some of the inputs. Our implementation is based on the differentiable inclusion-exclusion formulation of the hypervolume improvement (Daulton et al., 2020). As a by-product, the implementation also readily allows for the differentiable computation of the expected log hypervolume, instead of the log expected hypervolume, note the order, which can be preferable in certain applications of multi-objective optimization (Friedrich et al., 2011).
@@ 5 Empirical Results
We compare standard versions of analytic EI (EI) and constrained EI (CEI), Monte Carlo parallel EI (qEI), as well as Monte Carlo EHVI (qEHVI), in addition to other state-of-the-art baselines like lower-bound Max-Value Entropy Search (GIBBON) (Moss et al., 2021) and single- and multi-objective Joint Entropy Search (JES) (Hvarfner et al., 2022; Tu et al., 2022). All experiments are implemented using BoTorch (Balandat et al., 2020) and utilize multi-start optimization of the AF with scipy’s L-BFGS-B optimizer. In order to avoid conflating the effect of BoTorch’s default initialization strategy with those of our contributions, we use 16 initial points chosen uniformly at random from which to start the L-BFGS-B optimization. For a comparison with other initialization strategies, see Appendix D. We run multiple replicates and report mean and error bars of ±$\pm 2$ standard errors of the mean. Appendix D.1 contains additional details.
@@
In Figure 3, we compare performance on the Ackley and Michalewicz test functions (Surjanovic and Bingham, ). Notably, LogEI substantially outperforms EI on Ackley as the dimensionality increases. Ackley is a challenging multimodal function for which it is critical to trade off local exploitation with global exploration, a task made exceedingly difficult by the numerically vanishing gradients of EI in a large fraction of the search space. We see a similar albeit less pronounced behavior on Michalewicz, which reflects the fact that Michalewicz is a somewhat less challenging problem than Ackley.
@@
Figure 4 shows results on four engineering design problems with black box constraints that were also considered in (Eriksson and Poloczek, 2021). We apply the same bilog transform as the trust region-based SCBO method (Eriksson and Poloczek, 2021) to all constraints to make them easier to model with a GP. We see that LogCEI outperforms the naive CEI implementation and converges faster than SCBO. Similar to the unconstrained problems, the performance gains of LogCEI over CEI grow with increasing problem dimensionality and the number of constraints. Notably, we found that for some problems, LogCEI in fact improved upon some of the best results quoted in the original literature, while using three orders of magnitude fewer function evaluations, see Appendix D.7 for details.
@@
Figure 6 shows the performance of LogEI on three high-dimensional problems: the $6$-dimensional Hartmann function embedded in a $100$-dimensional space, a $100$-dimensional rover trajectory planning problem, and a $103$-dimensional SVM hyperparameter tuning problem. We use a $103$-dimensional version of the $388$-dimensional SVM problem considered by Eriksson and Jankowiak (2021), where the $100$ most important features were selected using Xgboost. Figure 6 shows that the optimization exhibits varying degrees of improvement from the inclusion of qLogEI, both when combined with SAASBO (Eriksson and Jankowiak, 2021) and a standard GP. In particular, qLogEI leads to significant improvements on the embedded Hartmann problem, even leading BO with the canonical GP to ultimately catch up with the SAAS-prior-equipped model. On the other hand, the differences on the SVM and Rover problems are not significant, see Section 6 for a discussion.
Figure 6: Best objective value as a function of number of function evaluations (iterations) on three high-dimensional problems, including Eriksson and Jankowiak (2021)’s SAAS prior.
@@
Figure 7 compares qLogEHVI and qEHVI on two multi-objective test problems with varying batch sizes, including the real-world-inspired cell network design for optimizing coverage and capacity (Dreifuerst et al., 2021). The results are consistent with our findings in the single-objective and constrained cases: qLogEHVI consistently outperforms qEHVI and even JES (Tu et al., 2022) for all batch sizes. Curiously, for the largest batch size and DTLZ2, qLogNEHVI’s improvement over the reference point (HV $>0$) occurs around three batches after the other methods, but dominates their performance in later batches. See Appendix D.5 for results on additional synthetic and real-world-inspired multi-objective problems such as the laser plasma acceleration optimization (Irshad et al., 2023a), and vehicle design optimization (Liao et al., 2008; Tanabe and Ishibuchi, 2020)
Figure 7: Batch optimization performance on two multi-objective problems, as measured by the hypervolume of the Pareto frontier across observed points. This plot includes JES (Tu et al., 2022). Similar to the single-objective case, the LogEI variant qLogEHVI significantly outperforms the baselines.

### v3

@@ 1 Introduction
Bayesian Optimization (BO) is a widely used and effective approach for sample-efficient optimization of expensive-to-evaluate black-box functions [26, 30], with applications ranging widely between aerospace engineering [50], biology and medicine [51], materials science [3], civil engineering [4], and machine learning hyperparameter optimization [68, 74]. BO leverages a probabilistic surrogate model in conjunction with an acquisition function to determine where to query the underlying objective function. Improvement-based acquisition functions, such as Expected Improvement (EI) and Probability of Improvement (PI), are among the earliest and most widely used acquisition functions for efficient global optimization of non-convex functions [60, 44]. EI has been extended to the constrained [31, 29], noisy [54], and multi-objective [20] setting, as well as their respective batch variants [79, 6, 13], and is a standard baseline in the BO literature [68, 26]. While much of the literature has focused on developing new sophisticated acquisition functions, subtle yet critical implementation details of foundational BO methods are often overlooked. Importantly, the performance of EI and its variants is inconsistent even for mathematically identical formulations and, as we show in this work, most often sub-optimal.
Although the problem of optimizing EI effectively has been discussed in various works, e.g. [26, 79, 33], prior focus has been on optimization algorithms and initialization strategies, rather than the fundamental issue of computing EI.
@@
    We extend the ideas behind analytical LogEI to other members of the EI family, including constrained EI (CEI), Expected Hypervolume Improvement (EHVI), as well as their respective batch variants for parallel BO, qEI and qEHVI, using smooth approximations of the acquisition utilities to obtain non-vanishing gradients. All of our methods are available as part of BoTorch [6].
@@
To increase the chance of finding the global optimum of non-convex functions, gradient-based optimization is typically performed from multiple starting points, which can help avoid getting stuck in local optima [72]. For improvement-based acquisition functions however, optimization becomes increasingly challenging as more data is collected and the likelihood of improving over the incumbent diminishes, see our theoretical results in Section 3 and the empirical illustration in Figure 1 and Appendix D.2. As a result, gradient-based optimization with multiple random starting points will eventually degenerate into random search when the gradients at the starting points are numerically zero. This problem is particularly acute in high dimensions and for objectives with a large range.
@@ 2.1 Gaussian Processes
Gaussian Processes (GP) [67] are the most widely used surrogates in BO, due to their high data efficiency and good uncertainty quantification. For our purposes, it suffices to consider a GP as a mapping that provides a multivariate Normal distribution over the outputs $f({\mathbf{x}})$ for any ${\mathbf{x}}$:
@@ 2.2 Improvement-based Acquisition Functions
For the fully-sequential ($q=1$), single-outcome ($M=1$) setting, “classic” EI [61] is defined as
@@
Constrained BO involves one or more black-box constraints and is typically formulated as finding $\max_{{\mathbf{x}}\in\mathbb{X}}f_{\textrm{true},1}({\mathbf{x}})$ such that $f_{\textrm{true},i}({\mathbf{x}})\leq 0$ for $i\in\{2,\dotsc,M\}$. Feasibility-weighting the improvement [29, 31] is a natural approach for this class of problems:
@@
In many settings, one may evaluate $f_{\textrm{true}}$ on $q>1$ candidates in parallel to increase throughput. The associated parallel or batch analogue of EI [32, 77] is given by
@@
Unlike EI, qEI does not admit a closed-form expression and is thus typically computed via Monte Carlo sampling, which also extends to non-Gaussian posterior distributions [77, 6]:
@@
due to observing a batch ${\mathbf{f}}({\mathbf{X}}):=[{\mathbf{f}}({\mathbf{x}}_{1}),\cdots,{\mathbf{f}}({\mathbf{x}}_{q})]$ of $q$ new observations. EHVI can be expressed in closed form if $q=1$ and the objectives are modeled with independent GPs [82], but Monte Carlo approximations are required for the general case (qEHVI) [13].
@@ 2.3 Optimizing Acquisition Functions
Optimizing an acquisition function (AF) is a challenging task that amounts to solving a non-convex optimization problem, to which multiple approaches and heuristics have been applied. These include gradient-free methods such as divided rectangles [43], evolutionary methods such as CMA-ES [34], first-order methods such as stochastic gradient ascent, see e.g., Wang et al. [77], Daulton et al. [15], and (quasi-)second order methods [26] such as L-BFGS-B [10]. Multi-start optimization is commonly employed with gradient-based methods to mitigate the risk of getting stuck in local minima. Initial points for optimization are selected via various heuristics with different levels of complexity, ranging from simple uniform random selection to BoTorch’s initialization heuristic, which selects initial points by performing Boltzmann sampling on a set of random points according to their acquisition function value [6]. See Appendix B for a more complete account of initialization strategies and optimization procedures used by popular implementations. We focus on gradient-based optimization as often leveraging gradients results in faster and more performant optimization [13].
Optimizing AFs for parallel BO that quantify the value of a batch of $q>1$ points is more challenging than optimizing their sequential counterparts due to the higher dimensionality of the optimization problem – $qd$ instead of $d$ – and the more challenging optimization surface. A common approach to simplify the problem is to use a sequential greedy strategy that greedily solves a sequence of single point selection problems. For $i=1,\dotsc,q$, candidate ${\mathbf{x}}_{i}$ is selected by optimizing the AF for $q=1$, conditional on the previously selected designs $\{{\mathbf{x}}_{1},...,{\mathbf{x}}_{i-1}\}$ and their unknown observations, e.g. by fantasizing the values at those designs [79]. For submodular AFs, including EI, PI, and EHVI, a sequential greedy strategy will attain a regret within a factor of $1/e$ compared to the joint optimum, and previous works have found that sequential greedy optimization yields improved BO performance compared to joint optimization [79, 13]. Herein, we find that our reformulations enable joint batch optimization to be competitive with the sequential greedy strategy, especially for larger batches.
@@ 2.4 Related Work
While there is a substantial body of work introducing a large variety of different AFs, much less focus has been on the question of how to effectively implement and optimize these AFs. Zhan and Xing [83] provide a comprehensive review of a large number of different variants of the EI family, but do not discuss any numerical or optimization challenges. Zhao et al. [84] propose combining a variety of different initialization strategies to select initial conditions for optimization of acquisition functions and show empirically that this improves optimization performance. However, they do not address any potential issues or degeneracies with the acquisition functions themselves. Recent works have considered effective gradient-based approaches for acquisition optimization. Wilson et al. [79] demonstrates how stochastic first-order methods can be leveraged for optimizing Monte Carlo acquisition functions. Balandat et al. [6] build on this work and put forth sample average approximations for MC acquisition functions that admit gradient-based optimization using deterministic higher-order optimizers such as L-BFGS-B.
Another line of work proposes to switch from BO to local optimization based on some stopping criterion to achieve faster local convergence, using either zeroth order [62] or gradient-based [59] optimization. While McLeod et al. [59] are also concerned with numerical issues, we emphasize that those issues arise due to ill-conditioned covariance matrices and are orthogonal to the numerical pathologies of improvement-based acquisition functions.
@@ 3 Theoretical Analysis of Expected Improvement’s Vanishing Gradients
For any given – and especially early – iteration, $\epsilon_{n}$ does not have to be small, as both the optimality gap and the maximal posterior standard deviation can be large initially. Note that under certain technical conditions on the kernel function and the asymptotic distribution of the training data $\mathcal{D}_{n}$, the maximum posterior variance vanishes guaranteeably as $n$ increases, see [52, Corollary 3.2]. On its own, Theorem 1 gives insight into the non-asymptotic behavior by exposing a dependence to the distribution of objective values $f$. In particular, if the set of inputs that give rise to high objective values ($\approx f^{*}$) is concentrated, $P(f({\mathbf{x}})<f^{*}-\epsilon)$ will decay very slowly as $\epsilon$ increases, thereby maintaining a lower bound on the probability of close to 1. As an example, this is the case for the Ackley function, especially as the dimensionality increases, which explains the behavior in Figure 1.
@@ 4.1 Analytic LogEI
where $c_{1}=\log(2\pi)/2$, and $c_{2}=\log(\pi/2)/2$, $\epsilon$ is the numerical precision, and log1mexp, erfcx are numerically stable implementations of $\log(1-\exp(z))$ and $\exp(z^{2})\text{erfc}(z)$, respectively, see App. A. Progenitors of Eq. (9) are found in SMAC 1.0 [37] and RoBO [48] which contain a log-transformed analytic EI implementation that is much improved, but can still exhibit instabilities as $z$ grows negative, see App. Fig. 10. To remedy similar instabilities, we put forth the third, asymptotic case in Eq. (9), ensuring numerical stability throughout, see App. A.2 for details. The asymptotically quadratic behavior of log_h becomes apparent in the last two cases, making the function particularly amenable to gradient-based optimization with significant practical implications for EI-based BO.
@@ 4.4 Monte Carlo Parallel LogEHVI
The numerical difficulties of qEHVI in (6) are similar to those of qEI, and the basic ingredients of smoothing and log-transformations still apply, but the details are significantly more complex since qEHVI uses many operations that have mathematically zero gradients with respect to some of the inputs. Our implementation is based on the differentiable inclusion-exclusion formulation of the hypervolume improvement [13]. As a by-product, the implementation also readily allows for the differentiable computation of the expected log hypervolume, instead of the log expected hypervolume, note the order, which can be preferable in certain applications of multi-objective optimization [28].
@@ 5 Empirical Results
We compare standard versions of analytic EI (EI) and constrained EI (CEI), Monte Carlo parallel EI (qEI), as well as Monte Carlo EHVI (qEHVI), in addition to other state-of-the-art baselines like lower-bound Max-Value Entropy Search (GIBBON) [63] and single- and multi-objective Joint Entropy Search (JES) [38, 73]. All experiments are implemented using BoTorch [6] and utilize multi-start optimization of the AF with scipy’s L-BFGS-B optimizer. In order to avoid conflating the effect of BoTorch’s default initialization strategy with those of our contributions, we use 16 initial points chosen uniformly at random from which to start the L-BFGS-B optimization. For a comparison with other initialization strategies, see Appendix D. We run multiple replicates and report mean and error bars of ±$\pm 2$ standard errors of the mean. Appendix D.1 contains additional details.
@@
In Figure 3, we compare performance on the Ackley and Michalewicz test functions [69]. Notably, LogEI substantially outperforms EI on Ackley as the dimensionality increases. Ackley is a challenging multimodal function for which it is critical to trade off local exploitation with global exploration, a task made exceedingly difficult by the numerically vanishing gradients of EI in a large fraction of the search space. We see a similar albeit less pronounced behavior on Michalewicz, which reflects the fact that Michalewicz is a somewhat less challenging problem than Ackley.
@@
Figure 4 shows results on four engineering design problems with black box constraints that were also considered in [22]. We apply the same bilog transform as the trust region-based SCBO method [22] to all constraints to make them easier to model with a GP. We see that LogCEI outperforms the naive CEI implementation and converges faster than SCBO. Similar to the unconstrained problems, the performance gains of LogCEI over CEI grow with increasing problem dimensionality and the number of constraints. Notably, we found that for some problems, LogCEI in fact improved upon some of the best results quoted in the original literature, while using three orders of magnitude fewer function evaluations, see Appendix D.7 for details.
@@
Figure 6 shows the performance of LogEI on three high-dimensional problems: the $6$-dimensional Hartmann function embedded in a $100$-dimensional space, a $100$-dimensional rover trajectory planning problem, and a $103$-dimensional SVM hyperparameter tuning problem. We use a $103$-dimensional version of the $388$-dimensional SVM problem considered by Eriksson and Jankowiak [21], where the $100$ most important features were selected using Xgboost. Figure 6 shows that the optimization exhibits varying degrees of improvement from the inclusion of qLogEI, both when combined with SAASBO [21] and a standard GP. In particular, qLogEI leads to significant improvements on the embedded Hartmann problem, even leading BO with the canonical GP to ultimately catch up with the SAAS-prior-equipped model. On the other hand, the differences on the SVM and Rover problems are not significant, see Section 6 for a discussion.
Figure 6: Best objective value as a function of number of function evaluations (iterations) on three high-dimensional problems, including Eriksson and Jankowiak [21]’s SAAS prior.
@@
Figure 7 compares qLogEHVI and qEHVI on two multi-objective test problems with varying batch sizes, including the real-world-inspired cell network design for optimizing coverage and capacity [19]. The results are consistent with our findings in the single-objective and constrained cases: qLogEHVI consistently outperforms qEHVI and even JES [73] for all batch sizes. Curiously, for the largest batch size and DTLZ2, qLogNEHVI’s improvement over the reference point (HV $>0$) occurs around three batches after the other methods, but dominates their performance in later batches. See Appendix D.5 for results on additional synthetic and real-world-inspired multi-objective problems such as the laser plasma acceleration optimization [40], and vehicle design optimization [56, 70]
Figure 7: Batch optimization performance on two multi-objective problems, as measured by the hypervolume of the Pareto frontier across observed points. This plot includes JES [73]. Similar to the single-objective case, the LogEI variant qLogEHVI significantly outperforms the baselines.